OpenStudy (anonymous):

Please help me find polar coordinates! /: These are rectangular coordinates given. (4,0)

OpenStudy (loser66):

it's (4, 0) but 4 is r =4 and 0 is \(\theta =0\) not as (4,0) which 4 is x =4 and 0 is y =0 ha!!! my explanation is ridiculous, not know whether you get me or not.

OpenStudy (loser66):

@jdoe0001 please help me explain it.

OpenStudy (jdoe0001):

hmm so... you have a rectangular coordinate set, (4, 0) x = 4, y = 0 and you'd want the polar coordinate set, which will be \(\bf (r, \theta)\)

OpenStudy (jdoe0001):

so, what's "r", or radius... well \(\bf r^2=x^2+y^2\quad \textit{from the pythagorean theorem}\\ \quad \\ x=4\quad y=0\quad thus\quad r^2=x^2+y^2\implies r^2=4^2+0^2\\ \quad \\ r=\sqrt{4^2+0^2}\) now, what's \(\theta\quad ? \ well, \(\bf tan(\theta)=\cfrac{y}{x}\implies \cfrac{0}{4}=0\qquad thus\\ \quad \\ tan(\theta)=0\implies tan^{-1}[tan(\theta)]=tan^{-1}(0)\implies \theta=tan^{-1}(0)\) once you find those 2 values, stick them as the coordinate ordered pair :)

OpenStudy (jdoe0001):

hmmm got a typo... but anyhow, \(\bf tan(\theta)=\cfrac{y}{x}\implies \cfrac{0}{4}=0\qquad thus\\ \quad \\ tan(\theta)=0\implies tan^{-1}[tan(\theta)]=tan^{-1}(0)\implies \theta=tan^{-1}(0) \)

OpenStudy (anonymous):

Thanks both of you! (: I also need help on two others... they are (3,4) and (2,-2) the more the help, the better but its okay if not (: Really appreciate it

OpenStudy (jdoe0001):

do the same thing, is pretty much just the same

OpenStudy (jdoe0001):

just keep in mind that \(\bf r=\sqrt{x^2+y^2}\qquad \theta=tan^{-1}\left(\cfrac{y}{x}\right)\)

OpenStudy (anonymous):

Thanks (:

OpenStudy (jdoe0001):

yw